# Ferromagnetic CrBr3-Induced Graphene Spintronics

### v3

@@ Abstract
Our proposed spin valve prototype features an advanced design with a two-dimensional graphene bilayer sandwiched between layers of ${\text{CrBr}}_{3}$ ferromagnetic insulators. In this configuration, proximity coupling plays a crucial role in determining the magnetization orientations of the graphene layers and significantly influences the in-plane conductivity of the ${\text{CrBr}}_{3}$ layers. In our study, we place the graphene bilayer between two ${\text{CrBr}}_{3}$ ferromagnetic insulator layers to create this spin valve structure. Utilizing density functional theory (DFT), we perform comprehensive calculations to investigate the electronic structure of this layered system. Our results reveal a substantial band gap of 56.87 meV at specific k-points, particularly pronounced in the antiparallel magnetization configuration. This finding marks a significant advancement in spintronics, highlighting the potential of our spin valve prototype to contribute to the development of cutting-edge electronic device technologies.
Introduction–Spin valves are crucial in spintronics, composed of layered ferromagnetic materials as conducting electrodes Castro Neto et al. (2009); Dieny et al. (1991). They are integral to modern electronics, leveraging giant magnetoresistance (MR) to switch between high and low resistance states Monsma et al. (1995); Park et al. (2011); Grünberg (2008); Lee et al. (2016). Understanding two-dimensional (2D) van der Waals (vdWs) systems has laid a robust foundation, offering finely tunable electronic properties in heterostructures Zhao et al. (2023); Li and He (2023); Paul et al. (2023); Flokstra et al. (2023); Wolf et al. (2021); Ojajärvi et al. (2022) with materials like h-BN, TMDCs (e.g., MoS2), graphene, and phosphorene Geim and Grigorieva (2013); Novoselov et al. (2016); Padilha et al. (2015); nat (2019); Han et al. (2014). In vdW spin valves, the tunable spin proximity effect (SPE) can induce a metal-insulator transition in graphene under strong magnetic fields, potentially leading to Quantum Hall ferromagnetism Kane and Mele (2005); Qiao et al. (2010); Zhang et al. (2012); Tse et al. (2011); Qiao et al. (2012); Liu et al. (2023); Zatko et al. (2023). Advances with monolayers of ferromagnetic insulator CrBr3 in vdWs devices have opened avenues for exploring spin-dependent transport phenomena Tombros et al. (2007); Song et al. (2018).
@@
Our study focuses on bilayer graphene, diverging from previous lateral spin valves that utilized monolayer graphene Tombros et al. (2007); Han et al. (2014). Monolayer graphene has limited contact area with ferromagnetic electrodes, unlike spin filter tunnel junctions Klein et al. (2018); Song et al. (2018); Wang et al. (2018). In contrast, our proposed van der Waals (vdW) spin valve integrates bilayer graphene between insulating magnetic layers Cardoso et al. (2018); Baibich et al. (1988); Binasch et al. (1989). This design offers advantages such as increased contact area between the graphene and magnetic materials, potentially enhancing spin transport properties. Moreover, bilayer graphene’s unique electronic structure and tunable properties enable precise control over spin-related phenomena. These attributes highlight the novelty and potential effectiveness of our vdW spin valve architecture.
@@
Computational Details–The geometry optimization and electronic structure calculations of the vdW heterostructure were conducted using the projector augmented wave (PAW) formalism Blöchl (1994) within density functional theory (DFT) Kresse and Furthmüller (1996). The Perdew-Burke-Ernzerhof (PBE) generalized-gradient approximation (GGA) was employed to describe the exchange and correlation functional Perdew et al. (1996). Specifically for the Cr 3d electrons, the GGA+U method was applied, with the effective on-site Coulomb interaction U set to 3.0 eV and exchange interaction J to 0.9 eV Liechtenstein et al. (1995). For the monolayer CrBr3, known for its accuracy but higher computational expense, the HSE06 hybrid functional Heyd et al. (2003) was adopted to achieve optimized geometry structures and electronic states. We design this heterostructure with a bilayer graphene (BLG) sheet surrounded by two monolayers of CrBr3. For the bilayer structure, we used a unit cell composed of 4 layers: CrBr3/BLG/CrBr3. In the superlattice arrangement, we adjusted the in-plane lattice parameter of graphene while varying the $c$ lattice parameter to optimize the interlayer distance (2.34 Å between CrBr3/BLG and 1.8 Å between BLG layers) and minimize the total energy. Each unit cell in the superlattice setup consisted of a 3$\times$3$\times$1 supercell, containing 14 carbon atoms per graphene layer, along with 2 chromium and 6 bromine atoms per CrBr3 layer. Consequently, the sandwiched graphene bilayer structure encompassed a total of 44 atoms. Our calculations revealed that the spin of Cr atoms is $S=3/2$, residing within the $t{2g}$ bands. A plane-wave cutoff energy of 550 eV and a vacuum space larger than 22 Å were implemented to prevent interactions between adjacent slabs. Grids of 9$\times$9$\times$1 and 27$\times$27$\times$1 Monkhorst-Pack k-point meshes were used for the vdW heterostructures and pure monolayer CrBr3, respectively, to integrate over the first Brillouin zone. Here, we consider both the parallel configurations showing ferromagnetic (FM) and antiferromagentic (AFM) orientations. The vdW correction was incorporated using the Grimme (DFT-D2) method Grimme (2006) in the heterostructure calculations. Electronic band structures were also computed using the Quantum ESPRESSO suite Giannozzi et al. (2009) to support and validate the computational findings.
@@
Results and Discussions–The atomic configuration is depicted in Fig.1(a-c), and the supercell architecture is illustrated in Fig.1(d-e). The interlayer distance between the graphene sheet and CrBr3 is 2.34 Å, while the interplanar distance between graphene sheets is fixed at 1.80 Å, characterized by van der Waals-type interlayer interactions. The total optimized energy is simulated as a function of interplanar distance to determine the optimal spacing for effective proximity coupling between the graphene sheet and CrBr3, incorporating dispersion corrections Monsma et al. (1995).
Dispersion correction has been integrated to accurately assess the bonding compatibility between the CrBr3 layer and the graphene sheet, ensuring precise determination of the total energy relative to the interplanar spacing. In constructing graphene-on-metal systems as heterostructures, it is crucial to comprehend and minimize junction resistance originating from the metal-graphene interface during geometry optimization, alongside implementing dispersion corrections for van der Waals interacting systems Park et al. (2011); Grünberg (2008); Han et al. (2014).
@@
These results highlight the potential of using a vdW magnetic system to explore quantum Hall effects, diverging from conventional ferromagnetic layer systems. Examining the band structure reveals fluctuations in the split-off energy gap between the ferromagnetic and antiferromagnetic states (illustrated in Fig.3(a)), along with the spin-dependent gap as a function of U across bias voltages from 0 to ±5 V/nm (shown in Fig.3(b)). This analysis demonstrates pronounced non-linearity under biasing conditions, indicative of interlayer charge polarization. Moreover, the mixed (both linear and nonlinear) behavior prompts further investigation into charge transmission spectra to understand the proximity effect in heterostructures and the role of interlayer polarization for potential device applications (as depicted in the inset heterostructure in Fig. 2). We now explore the origin of gap formation in the graphene bilayer within an antiferromagnetic aligned vdW spin valve. In this setup, the Hamiltonian for each spin channel mirrors the model of a graphene bilayer under an off-plane electric field, which is known to induce a band gap in the structure Ohta et al. (2006); McCann (2006). Notably, within the spin valve configuration, the effective electric field exhibits opposite signs for different spin orientations, $E_{\rm eff}\propto\sigma\Delta$. The spin projection of the antiferromagnetic bands in the top and bottom layers, as depicted in Fig. 3(b), distinctly manifests a spin dipole effect: for a given spin direction, there arises a charge imbalance induced by the exchange interaction with the magnetic layers, precisely offset by the opposite spin direction.

### v4

@@ Abstract
Introduction–Spin valves are crucial in spintronics, composed of layered ferromagnetic materials as conducting electrodes RevModPhys.81.109; PhysRevB.43.1297. They are integral to modern electronics, leveraging giant magnetoresistance (MR) to switch between high and low resistance states monsma1995perpendicular; park2011aspinvalve; grunberg2008from; acs.nanolett.6b03052. Understanding two-dimensional (2D) van der Waals (vdWs) systems has laid a robust foundation, offering finely tunable electronic properties in heterostructures adma.202209113; acs.jpclett.3c03000; sciadv.abn1401; Flokstra2023; PhysRevLett.126.056803; PhysRevLett.128.167701 with materials like h-BN, TMDCs (e.g., MoS2), graphene, and phosphorene Geim2013; Novoselovaac9439; PhysRevLett.114.066803; nat2019years; han2014graphene. In vdW spin valves, the tunable spin proximity effect (SPE) can induce a metal-insulator transition in graphene under strong magnetic fields, potentially leading to Quantum Hall ferromagnetism PhysRevLett.95.146802; PhysRevB.82.161414; PhysRevLett.108.056802; PhysRevB.83.155447; PhysRevB.85.115439; acs.nanolett.3c02489; acs.nanolett.2c03113. Advances with monolayers of ferromagnetic insulator CrBr3 in vdWs devices have opened avenues for exploring spin-dependent transport phenomena Tombros2007; Song1214.
Recently, a van der Waals (vdW) spin valve was demonstrated using bilayer graphene (BLG) between CrI3 ferromagnetic insulators PhysRevLett.121.067701. In our study, we aim to explore a similar configuration but replacing CrI3 with CrBr3, which has a slightly lower Curie temperature (TC) of 33K PhysRevB.56.719 compared to CrI3 (TC of 45K) Huang2017; PhysRevLett.121.067701. CrBr3 is notable for its heavy Cr atoms, suggesting potential for strong spin-orbit coupling (SOC) effects in graphene PhysRevB.84.195444; PhysRevLett.109.055502; acsami.4c01034. Its hexagonal crystal structure aligns well with graphene’s lattice, with a modest lattice mismatch of approximately 4.5 $\%$. This structural compatibility makes CrBr3 promising for vdW heterostructures. The layered architecture of CrBr3, bound by vdW forces, allows easy exfoliation into 2D monolayers while maintaining ferromagnetic insulating properties PhysRevLett.110.046603; Hallal_2017; acs.nanolett.9b00553, enhancing the versatility of vdW spin valves. In CrBr3 bilayers, weak interlayer coupling is conducive to controlling layer magnetizations independently, crucial for spintronic applications PhysRevB.84.195444; PhysRevLett.109.055502. This exploration opens new avenues for studying spin-dependent phenomena in 2D vdW heterostructures.
Our study focuses on bilayer graphene, diverging from previous lateral spin valves that utilized monolayer graphene Tombros2007; han2014graphene. Monolayer graphene has limited contact area with ferromagnetic electrodes, unlike spin filter tunnel junctions Klein1218; Song1214; Wang2018. In contrast, our proposed van der Waals (vdW) spin valve integrates bilayer graphene between insulating magnetic layers PhysRevLett.121.067701; PhysRevLett.61.2472; PhysRevB.39.4828. This design offers advantages such as increased contact area between the graphene and magnetic materials, potentially enhancing spin transport properties. Moreover, bilayer graphene’s unique electronic structure and tunable properties enable precise control over spin-related phenomena. These attributes highlight the novelty and potential effectiveness of our vdW spin valve architecture.
@@
Computational Details–The geometry optimization and electronic structure calculations of the vdW heterostructure were conducted using the projector augmented wave (PAW) formalism PhysRevB.50.17953 within density functional theory (DFT) PhysRevB.54.11169. The Perdew-Burke-Ernzerhof (PBE) generalized-gradient approximation (GGA) was employed to describe the exchange and correlation functional PhysRevLett.77.3865. Specifically for the Cr 3d electrons, the GGA+U method was applied, with the effective on-site Coulomb interaction U set to 3.0 eV and exchange interaction J to 0.9 eV PhysRevB.52.R5467. For the monolayer CrBr3, known for its accuracy but higher computational expense, the HSE06 hybrid functional 10.1063/1.1564060 was adopted to achieve optimized geometry structures and electronic states. We design this heterostructure with a bilayer graphene (BLG) sheet surrounded by two monolayers of CrBr3. For the bilayer structure, we used a unit cell composed of 4 layers: CrBr3/BLG/CrBr3. In the superlattice arrangement, we adjusted the in-plane lattice parameter of graphene while varying the $c$ lattice parameter to optimize the interlayer distance (2.34 Å between CrBr3/BLG and 1.8 Å between BLG layers) and minimize the total energy. Each unit cell in the superlattice setup consisted of a $\times3$$\times$1 supercell, containing 14 carbon atoms per graphene layer, along with 2 chromium and 6 bromine atoms per CrBr3 layer. Consequently, the sandwiched graphene bilayer structure encompassed a total of 44 atoms. Our calculations revealed that the spin of Cr atoms is $S=3/2$, residing within the t2gt{2g} bands. A plane-wave cutoff energy of 550 eV and a vacuum space larger than 22 Å were implemented to prevent interactions between adjacent slabs. Grids of $\times9$$\times$1 and $\times27$$\times$1 Monkhorst-Pack k-point meshes were used for the vdW heterostructures and pure monolayer CrBr3, respectively, to integrate over the first Brillouin zone. Here, we consider both the parallel configurations showing ferromagnetic (FM) and antiferromagentic (AFM) orientations. The vdW correction was incorporated using the Grimme (DFT-D2) method 10.1002.20495 in the heterostructure calculations. Electronic band structures were also computed using the Quantum ESPRESSO suite Giannozzi_2009 to support and validate the computational findings.
@@
Results and Discussions–The atomic configuration is depicted in Fig.1(a-c), and the supercell architecture is illustrated in Fig.1(d-e). The interlayer distance between the graphene sheet and CrBr3 is 2.34 Å, while the interplanar distance between graphene sheets is fixed at 1.80 Å, characterized by van der Waals-type interlayer interactions. The total optimized energy is simulated as a function of interplanar distance to determine the optimal spacing for effective proximity coupling between the graphene sheet and CrBr3, incorporating dispersion corrections monsma1995perpendicular.
Dispersion correction has been integrated to accurately assess the bonding compatibility between the CrBr3 layer and the graphene sheet, ensuring precise determination of the total energy relative to the interplanar spacing. In constructing graphene-on-metal systems as heterostructures, it is crucial to comprehend and minimize junction resistance originating from the metal-graphene interface during geometry optimization, alongside implementing dispersion corrections for van der Waals interacting systems park2011aspinvalve; grunberg2008from; han2014graphene.
@@
These results highlight the potential of using a vdW magnetic system to explore quantum Hall effects, diverging from conventional ferromagnetic layer systems. Examining the band structure reveals fluctuations in the split-off energy gap between the ferromagnetic and antiferromagnetic states (illustrated in Fig.3(a)), along with the spin-dependent gap as a function of U across bias voltages from 0 to ±5 V/nm (shown in Fig.3(b)). This analysis demonstrates pronounced non-linearity under biasing conditions, indicative of interlayer charge polarization. Moreover, the mixed (both linear and nonlinear) behavior prompts further investigation into charge transmission spectra to understand the proximity effect in heterostructures and the role of interlayer polarization for potential device applications (as depicted in the inset heterostructure in Fig. 2). We now explore the origin of gap formation in the graphene bilayer within an antiferromagnetic aligned vdW spin valve. In this setup, the Hamiltonian for each spin channel mirrors the model of a graphene bilayer under an off-plane electric field, which is known to induce a band gap in the structure Ohta951; PhysRevB.74.161403. Notably, within the spin valve configuration, the effective electric field exhibits opposite signs for different spin orientations, $E_{\rm eff}\propto\sigma\Delta$. The spin projection of the antiferromagnetic bands in the top and bottom layers, as depicted in Fig. 3(b), distinctly manifests a spin dipole effect: for a given spin direction, there arises a charge imbalance induced by the exchange interaction with the magnetic layers, precisely offset by the opposite spin direction.